OpenStudy (anonymous):

Which is the factored form of x3 - 10x2 + 27x -18 = 0? (Points : 4) (x + 1)(x- 3)(x -6) = 0 (x + 1)(x -2)(x + 9) = 0 (x -1)(x- 3)(x- 6) = 0 (x- 1)(x- 2)(x + 9) = 0

OpenStudy (welshfella):

you missed out some signs

OpenStudy (freckles):

\[x^3 10x^2+27x18=0\]?

OpenStudy (freckles):

I don't believe the left hand side has any of the factors of the left hand side of the equations you have given as choices

OpenStudy (anonymous):

well it was what it was givin to me

OpenStudy (freckles):

no but it looks like x^3-10x^2+27x-18=0 was given to you

OpenStudy (anonymous):

yea

OpenStudy (freckles):

see if you can a zero try some factors of -18 and see which gives you 0

OpenStudy (welshfella):

Hint; if one factor is (x -1) then f(1) = 0

OpenStudy (anonymous):

it would be either the first answer or last answer

OpenStudy (freckles):

if it is the first answer then you are saying -1,3,6 are the zeros if you are saying it is the last answer then you are saying 1,2,-9 are the zeros plug these into the initial function to see which outputs zeros for all those x's

OpenStudy (anonymous):

what numbers do I use to plug into them

OpenStudy (anonymous):

no wait it is the 3rd answer